# Unraveling the Smoothness Properties of Diffusion Models: A Gaussian Mixture Perspective

### v1

@@
Jiuxiang Gu jigu@adobe.com. Adobe Research.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song zsong@adobe.com. Adobe Research.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
@@ 1 Introduction
However, despite these technological strides, a critical gap persists in the theoretical understanding of diffusion processes, especially concerning the Lipschitz continuity and second momentum properties of these models. Many existing studies ([CCL+22, LLT22, CLL23, CCL+23] and many more) make simplifying assumptions about these key smoothness properties but lack rigorous formulation or comprehensive analysis. This paper aims to bridge this theoretical gap by providing a detailed examination of the Lipschitz and second momentum characteristics. To make the data distribution analyzable, we consider the mixture of Gaussian data distribution, which is a universal approximation (Fact 1.1) for any smooth density, such as complex image and video data distributions.
@@
Our contribution:
Other studies of the mixture of Gaussian under diffusion. Recently, there has been a rich line of work that studies a mixture of Gaussian data distribution under the diffusion process, which shares a similar popular setting. However, none focus on the Lipschitz smoothness and second momentum property as ours. We provide a short summary here for convenience. [WCL+24] analyses the effect of diffusion guidance and provides theoretical insights on the instillation of task-specific information into the score function and conditional generation, e.g., text-to-image, by studying the mixture of Gaussian data distribution as a case study. [GLB+24] propose a new SDE-solver for diffusion models under a mixture of Gaussian data. [SCK23, GKL24] learn mixtures of Gaussian data using the $\mathsf{DDPM}$ objective and gives bound for sample complexity by assuming all covariance matrices are identity but does not use Lipschitz, while our work does not need identity assumptions. [CKS24] mainly focuses on solving $k$-mixture of Gaussian in diffusion model by leveraging the property of the mixture of Gaussian and polynomial approximation methods and gives bound for sample complexity by assuming the covariance matrices have bounded condition number and that the mean vectors and covariance matrices lie in a bounded ball.
@@ 2 Related Work
Diffusion models and score-based generative models. [HJA20] introduced the concept of denoising diffusion probabilistic models ($\mathsf{DDPM}$), which learn to reverse a gradual noising process to generate samples. Then, [SSDK+20] used Stochastic Differential Equations (SDE) to build the diffusion model and explored the equivalence between score-based generative models (SGMs) and diffusion models, generalizing the diffusion model to continuous time. There is a line of works [SDWMG15] studying the connection between diffusion models and non-equilibrium thermodynamics, providing a theoretical foundation for these models. Another line of work has focused on improving the efficiency and quality of diffusion models [NRW21]. Particularly, [SE19] leverages score matching to train diffusion models more efficiently; [ND21] improved architectures and training techniques for diffusion models. Diffusion models have been applied to various domains beyond image generation, achieving impressive results, e.g., text-to-image synthesis [CZZ+20], audio synthesis [KPH+20], image super-resolution [SHC+22], and so on.
Mixture of Gaussian. Mixtures of Gaussian are among the most fundamental and widely used statistical models [Das99, FOS08, MV10, BS10, CDSS13, SOAJ14, DK14, DKK+19, ADLS17, LS17, ABDH+18, DK20, BK20, DHKK20, SCL+22, BDJ+22, XSW+23, BS23, SMF+24] and studied in neural networks learning [RGKZ21, SWL21, SWL24]. Recently, they have also been widely studied in diffusion models [SCK23, WCL+24, GLB+24, GKL24, CKS24] (see detailed discussion in Section 1).
Lipschitz and second momentum in score estimation. There is a line of work using bounded Lipschitz and second momentum as their key smoothness assumptions for the whole diffusion process without giving any concrete formulation [CCL+22, KFL22a, LLT22, LKB+23a, CLL23, CHZW23, CCL+23, CDD23, BDD23, ZHF+24, KLL+24], while our work gives a “close-form” formulation of Lipschitz and second momentum. There is another rich line of work studying how to train the diffusion models to have a better theoretical guarantee [SE19, SE20, SK21, SGSE20, SDME21, KFL22b, SDCS23, LKB+23b, CDD23, CHZW23, SCK23, YFZ+23, GLL+24, GKL24, CCL+23, GLB+24, CKS24, HRX24] and many more.
Roadmap.
In Section 3, we provide the notations we use, several definitions and lemmas related to $k$ mixtures of Gaussian. In Section 4, we provide the preliminary knowledge on score-based generative models (SGMs) and diffusion models. In Section 5, we present our main results. In Section 6, we conclude the paper.
@@ 3 Lipschitz and Second Momentum of Mixture of Gaussian
In this section, we discuss the notations used, several key definitions for $k$ mixtures of Gaussian distributions, and lemmas concerning the Lipschitz continuity and second momentum of these mixtures. We begin by presenting the notations that are used throughout the paper. Notations. For two vectors $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$, i.e., $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. We use $\Pr[]$ to denote the probability, we use $\operatorname*{{\mathbb{E}}}[]$ to denote the expectation. We use $e_{i}$ to denote a vector where only $i$-th coordinate is $1$, and other entries are 00. For each $a,b\in\mathbb{R}^{n}$, we use $a\circ b\in\mathbb{R}^{n}$ to denote the vector where $i$-th entry is $(a\circ b)_{i}=a_{i}b_{i}$ for all $i\in[n]$, and this is the Hardamard product. We use ${\bf 1}_{n}$ to denote a length-$n$ where all the entries are ones. We use $x_{i,j}$ to denote the $j$-th coordinate of $x_{i}\in\mathbb{R}^{n}$. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e. $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$. For $k>n$, for any matrix $A\in\mathbb{R}^{k\times n}$, we use $\|A\|$ to denote the spectral norm of $A$, i.e. $\|A\|:=\sup_{x\in\mathbb{R}^{n}}\|Ax\|_{2}/\|x\|_{2}$. We use $\sigma_{\min}(A),\sigma_{\max}(A)$ to denote the minimum/maximum singular value of matrix $A$. For a square matrix $A$, we use $\operatorname{tr}[A]$ to denote the trace of $A$, i.e., $\operatorname{tr}[A]=\sum_{i=1}^{n}A_{i,i}$. We use $\det(A)$ to denote the determinant of matrix $A$. We use $f*g$ to denote the convolution of 2 functions $f,g$. In addition to $O(\cdot)$ notation, for two functions $f,g$, we use the shorthand $f\lesssim g$ (resp. $\gtrsim$) to indicate that $f\leq Cg$ (resp. $\geq$) for an absolute constant $C$.
@@
$p_{t}(x):=\sum_{i=1}^{k}\frac{\alpha_{i}(t)}{(2\pi)^{d/2}\det(\Sigma_{i}(t))^{1/2}}\exp(-\frac{1}{2}(x-\mu_{i}(t))^{\top}\Sigma_{i}(t)^{-1}(x-\mu_{i}(t))).$
Then, the following lemma shows that the linear combination between $k$ mixtures of Gaussian and a single standard Gaussian is still a $k$ mixtures of Gaussian. The proof is in Appendix H.
Lemma 3.2 (Informal version of Lemma H.1).
@@
Main properties. Thus, we only need to analyze the property, i.e., Lipschitz constant and second momentum, of $k$ mixtures of Gaussian, to have a clear guarantee for the whole diffusion process. In the following two lemmas, we will give concrete bounds of the Lipschitz constant and second momentum of $k$ mixtures of Gaussian. Both proofs can be found in Appendix H.
Lemma 3.3 (Lipschitz, informal version of Lemma H.3).
@@
$L=\frac{1}{\sigma_{\min(p_{t})}}+\frac{2R^{2}}{\gamma^{2}\sigma_{\min(p_{t})}^{2}}\cdot(\frac{1}{(2\pi)^{d}\det_{\min(p_{t})}}+\frac{1}{(2\pi)^{d/2}\det_{\min(p_{t})}^{1/2}})\cdot\exp(-\frac{\beta^{2}}{2\sigma_{\max(p_{t})}}).$
@@
We believe our results in Lemma 3.3 is non-trivial. Note that, in practice, the $k$ will be super large for complicated data distribution, e.g., millions of Gaussian components for an image distribution. Many studies need to overcome the large $k$ hardness in learning Gaussian mixture models [BDJ+22, BS23, RGKZ21, SWL24], while our Lipschitz upper-bound is independent with $k$.
Lemma 3.5 (Second momentum, informal version of Lemma H.2).
@@
The proof idea of Lemma 3.3 and Lemma 3.5 is that we first consider the case when $k=1$ in Appendix E and then we extend to $2$ mixtures of Gaussian setting in Appendix F. Finally, we can generalize to $k$ setting in Appendix G and summarize in Appendix H.
@@ 4.1 Background on score based model and diffusion model
$\mathrm{d}\widetilde{x}_{t}=(-f(\widetilde{x}_{t},T-t)+g(T-t)^{2}\nabla\log p_{T-t}(\widetilde{x}_{t}))\mathrm{d}t+g(T-t)\mathrm{d}w_{t}.$ (4)
@@
$\mathrm{d}y_{t}=(-f(y_{t},T-t)+g(T-t)^{2}s_{T-t}(y_{t}))\mathrm{d}t+g(T-t)\,\mathrm{d}w_{t},~{}~{}y_{0}\sim p_{T},~{}~{}0\leq t\leq T.$ (5)
@@
The forward SDE of OU process ($f(x_{t},t)=-x_{t}$, and $g(t)\equiv\sqrt{2}$) is:
$\mathrm{d}x_{t}=-x_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}w_{t}.$
The corresponding discrete Markov chain form of OU process is given by:
$x_{t}=e^{-t}x_{0}+\sqrt{1-e^{-2t}}z,$
where we can see that $a_{t}=e^{-t}$, $b_{t}=\sqrt{1-e^{-2t}}$ in Eq. (2).
@@
$\mathrm{d}y_{t}=(x_{t}+2s_{T-t}(y_{t}))\mathrm{d}t+\sqrt{2}\mathrm{d}w_{t}$ (6)
@@ 4.2 Definitions of different solvers
where $\widehat{y}_{T-\delta}\in\mathbb{R}^{d}$ is the output, $T\in\mathbb{R}^{+}$ is the total time, and $s$ is the score estimates. The Euler-Maruyama scheme is, (Equation 5 in [CLL23]),
$\mathrm{d}\widehat{y}_{t}=[\widehat{y}_{t^{\prime}_{k}}+2s_{T-t^{\prime}_{k}}(\widehat{y}_{t^{\prime}_{k}})]\mathrm{d}t+\sqrt{2}\mathrm{d}w_{t},\quad t\in[t_{k}^{\prime},t_{k+1}^{\prime}].$ (7)
@@
where $\widehat{y}_{T-\delta}\in\mathbb{R}^{d}$ is the output, $T\in\mathbb{R}^{+}$ is the total time, and $s$ is the score estimates. The Exponential Integrator scheme is, (Equation 6 in [CLL23]),
$\mathrm{d}\widehat{y}_{t}=[\widehat{y}_{t}+2s_{T-t^{\prime}_{k}}(\widehat{y}_{t^{\prime}_{k}})]\mathrm{d}t+\sqrt{2}\mathrm{d}w_{t},\quad t\in[t_{k}^{\prime},t_{k+1}^{\prime}].$ (8)
@@
We define $\mathsf{DPOM}$ (Diffusion Predictor with Overdamped Modeling) as Algorithm 1 in [CCL+23] satisfying
$\widehat{y}_{T-\delta}=\mathsf{DPOM}(T,h_{\mathrm{pred}},h_{\mathrm{corr}},s),$
where $\delta$ is the early stopping parameter, $\widehat{y}_{T-\delta}\in\mathbb{R}^{d}$ is the output, $T\in\mathbb{R}^{+}$ is the total steps, $h_{\mathrm{pred}}$ is the predictor step size, $h_{\mathrm{corr}}$ is the corrector step size, (see detailed definition in Algorithm 1 of [CCL+23]), and $s$ is the score estimates.
@@
We define $\mathsf{DPUM}$ (Diffusion Predictor with Underdamped Modeling) as Algorithm 2 in [CCL+23], satisfying
$\widehat{y}_{T-\delta}=\mathsf{DPUM}(T,h_{\mathrm{pred}},h_{\mathrm{corr}},s),$
where $\delta$ is the early stopping parameter, $\widehat{y}_{T-\delta}\in\mathbb{R}^{d}$ is the output, $T\in\mathbb{R}^{+}$ is the total steps, $h_{\mathrm{pred}}$ is the predictor step size, $h_{\mathrm{corr}}$ is the corrector step size, (see detailed definition in Algorithm 2 of [CCL+23]), and $s$ is the score estimates.

### v2

@@
Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
@@ 1 Introduction
However, despite these technological strides, a critical gap persists in the theoretical understanding of diffusion processes, especially concerning the Lipschitz continuity and second momentum properties of these models. Many existing studies ([DB22, CCL+22, LLT22, CLL23, CCL+23] and many more) make simplifying assumptions about these key smoothness properties but lack rigorous formulation or comprehensive analysis. This paper aims to bridge this theoretical gap by providing a detailed examination of the Lipschitz and second momentum characteristics. To make the data distribution analyzable, we consider the mixture of Gaussian data distribution, which is a universal approximation (Fact 1.1) for any smooth density, such as complex image and video data distributions.
@@
Our contribution:
Other studies of the mixture of Gaussian under diffusion.
Recently, there has been a rich line of work that studies a mixture of Gaussian data distribution under the diffusion process, which shares a similar popular setting. However, none focus on the Lipschitz smoothness and second momentum property as ours. We provide a short summary here for convenience. [WCL+24] analyses the effect of diffusion guidance and provides theoretical insights on the instillation of task-specific information into the score function and conditional generation, e.g., text-to-image, by studying the mixture of Gaussian data distribution as a case study. [GLB+24] propose a new SDE-solver for diffusion models under a mixture of Gaussian data. [SCK23, GKL24] learn mixtures of Gaussian data using the $\mathsf{DDPM}$ objective and gives bound for sample complexity by assuming all covariance matrices are identity but does not use Lipschitz, while our work does not need identity assumptions. [CKS24] mainly focuses on solving $k$-mixture of Gaussian in diffusion model by leveraging the property of the mixture of Gaussian and polynomial approximation methods and gives bound for sample complexity by assuming the covariance matrices have bounded condition number and that the mean vectors and covariance matrices lie in a bounded ball.
@@ 2 Related Work
Diffusion models and score-based generative models. [HJA20] introduced the concept of denoising diffusion probabilistic models ($\mathsf{DDPM}$), which learn to reverse a gradual noising process to generate samples and have been applied to many area [WSD+23, WCZ+23, WXZ+24]. Then, [SSDK+20] used Stochastic Differential Equations (SDE) to build the diffusion model and explored the equivalence between score-based generative models (SGMs) and diffusion models, generalizing the diffusion model to continuous time. There is a line of works [SDWMG15] studying the connection between diffusion models and non-equilibrium thermodynamics, providing a theoretical foundation for these models. Another line of work has focused on improving the efficiency and quality of diffusion models [NRW21]. Particularly, [SE19] leverages score matching to train diffusion models more efficiently; [ND21] improved architectures and training techniques for diffusion models. Diffusion models have been applied to various domains beyond image generation, achieving impressive results, e.g., text-to-image synthesis [CZZ+20], audio synthesis [KPH+20], image super-resolution [SHC+22], and so on.
Mixture of Gaussian. Mixtures of Gaussian are among the most fundamental and widely used statistical models [Das99, FOS08, MV10, BS10, CDSS13, SOAJ14, DK14, DKK+19, ADLS17, LS17, ABDH+18, DK20, BK20, DHKK20, SCL+22, BDJ+22, XSW+23, BS23, SMF+24] and studied in neural networks learning [RGKZ21, SWL21, LSSS24, SWL24]. Recently, they have also been widely studied in diffusion models [SCK23, WCL+24, GLB+24, GKL24, CKS24] (see detailed discussion in Section 1).
Lipschitz and second momentum in score estimation. There is a line of work using bounded Lipschitz and second momentum as their key smoothness assumptions for the whole diffusion process without giving any concrete formulation [CCL+22, KFL22, LLT22, LKB+23, CLL23, CHZW23, CCL+23, CDD23, BDD23, ZHF+24, KLL+24], while our work gives a “close-form” formulation of Lipschitz and second momentum. There is another rich line of work studying how to train the diffusion models to have a better theoretical guarantee [SE19, SE20, SK21, SGSE20, SDME21, SDCS23, LKB+23, CDD23, RHX+23, CHZW23, SCK23, YFZ+23, LLSS24, GKL24, CCL+23, GLB+24, CKS24, HRX24, HWSL24] and many more.
Roadmap. In Section 3, we provide the notation we use, several definitions, and lemmas related to $k$ mixtures of Gaussian. In Section 4, we provide preliminary knowledge on score-based generative models (SGMs) and diffusion models. Section 6 provides the tools that we use from previous papers. In Section 5, we present our main results. In Section 7, we conclude the paper.
@@ 3 Lipschitz and Second Momentum of Mixture of Gaussian
In this section, we discuss the notations used, several key definitions for $k$ mixtures of Gaussian distributions, and lemmas concerning the Lipschitz continuity and second momentum of these mixtures. We begin by presenting the notations that are used throughout the paper. Notations. For two vectors $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$, i.e., $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. We use $\Pr[]$ to denote the probability, we use $\operatorname*{{\mathbb{E}}}[]$ to denote the expectation. We use $e_{i}$ to denote a vector where only $i$-th coordinate is $1$, and other entries are 00. For each $a,b\in\mathbb{R}^{n}$, we use $a\circ b\in\mathbb{R}^{n}$ to denote the vector where $i$-th entry is $(a\circ b)_{i}=a_{i}b_{i}$ for all $i\in[n]$, and this is the Hardamard product. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $x_{i,j}$ to denote the $j$-th coordinate of $x_{i}\in\mathbb{R}^{n}$. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e. $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$. For $k>n$, for any matrix $A\in\mathbb{R}^{k\times n}$, we use $\|A\|$ to denote the spectral norm of $A$, i.e. $\|A\|:=\sup_{x\in\mathbb{R}^{n}}\|Ax\|_{2}/\|x\|_{2}$. We use $\sigma_{\min}(A),\sigma_{\max}(A)$ to denote the minimum/maximum singular value of matrix $A$. For a square matrix $A$, we use $\operatorname{tr}[A]$ to denote the trace of $A$, i.e., $\operatorname{tr}[A]=\sum_{i=1}^{n}A_{i,i}$. We use $\det(A)$ to denote the determinant of matrix $A$. We use $f*g$ to denote the convolution of 2 functions $f,g$. In addition to $O(\cdot)$ notation, for two functions $f,g$, we use the shorthand $f\lesssim g$ (resp. $\gtrsim$) to indicate that $f\leq Cg$ (resp. $\geq$) for an absolute constant $C$.
@@
$p_{t}(x):=$ $~{}\sum_{i=1}^{k}\frac{\alpha_{i}(t)}{(2\pi)^{d/2}\det(\Sigma_{i}(t))^{1/2}}\cdot\exp(-\frac{1}{2}(x-\mu_{i}(t))^{\top}\Sigma_{i}(t)^{-1}(x-\mu_{i}(t))).$
Then, the following lemma shows that the linear combination between $k$ mixtures of Gaussian and a single standard Gaussian is still a $k$ mixtures of Gaussian. The proof is in Appendix G.
Lemma 3.2 (Informal version of Lemma G.1).
@@
$p(x):=$ $~{}\sum_{i=1}^{k}\frac{\alpha_{i}}{(2\pi)^{d/2}\det(\Sigma_{i})^{1/2}}\cdot\exp(-\frac{1}{2}(x-\mu_{i})^{\top}\Sigma_{i}^{-1}(x-\mu_{i}))$
@@
$\widetilde{p}(x):=$ $~{}\sum_{i=1}^{k}\frac{\alpha_{i}}{(2\pi)^{d/2}\det(\widetilde{\Sigma}_{i})^{1/2}}\cdot\exp(-\frac{1}{2}(x-\widetilde{\mu}_{i})^{\top}\widetilde{\Sigma}_{i}^{-1}(x-\widetilde{\mu}_{i})),$
@@
Main properties. Thus, we only need to analyze the property, i.e., Lipschitz constant and second momentum, of $k$ mixtures of Gaussian, to have a clear guarantee for the whole diffusion process. In the following two lemmas, we will give concrete bounds of the Lipschitz constant and second momentum of $k$ mixtures of Gaussian. Both proofs can be found in Appendix G.
Lemma 3.3 (Lipschitz, informal version of Lemma G.3).
@@
$L=$ $~{}\frac{1}{\sigma_{\min(p_{t})}}+\frac{2R^{2}}{\gamma^{2}\sigma_{\min(p_{t})}^{2}}\cdot\exp(-\frac{\beta^{2}}{2\sigma_{\max(p_{t})}})\cdot(\frac{1}{(2\pi)^{d}\det_{\min(p_{t})}}+\frac{1}{(2\pi)^{d/2}\det_{\min(p_{t})}^{1/2}}).$
@@
We believe our results in Lemma 3.3 is non-trivial. First, the Lipschitz bound depends inversely exponential on the dimension $d$. If $d$ grows lager, the Lipschitz constant $L$ becomes much smaller. Second, in practice, the $k$ will be super large for complicated data distribution, e.g., millions of Gaussian components for an image distribution. Many studies need to overcome the large $k$ hardness in learning Gaussian mixture models [BDJ+22, BS23, RGKZ21, SWL24], while our Lipschitz upper-bound is independent with $k$.
Lemma 3.5 (Second momentum, informal version of Lemma G.2).
@@
The proof idea of Lemma 3.3 and Lemma 3.5 is that we first consider the case when $k=1$ in Appendix D and then we extend to $2$ mixtures of Gaussian setting in Appendix E. Finally, we can generalize to $k$ setting in Appendix F and summarize it in Appendix G.
@@ 4.1 Background on score based model and diffusion model
$\mathrm{d}\widetilde{x}_{t}=$ $~{}(-f(\widetilde{x}_{t},T-t)+g(T-t)^{2}\nabla\log p_{T-t}(\widetilde{x}_{t}))\mathrm{d}t+g(T-t)\mathrm{d}w_{t}.$ (4)
@@
$\mathrm{d}y_{t}=$ $~{}(-f(y_{t},T-t)+g(T-t)^{2}s_{T-t}(y_{t}))\mathrm{d}t+g(T-t)\,\mathrm{d}w_{t},~{}~{}y_{0}\sim p_{T},~{}~{}0\leq t\leq T.$ (5)
@@
The forward SDE of OU process ($f(x_{t},t)=-x_{t}$, and $g(t)\equiv\sqrt{2}$) is: $\mathrm{d}x_{t}=-x_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}w_{t}.$ The corresponding discrete Markov chain form of OU process is given by: $x_{t}=e^{-t}x_{0}+\sqrt{1-e^{-2t}}z,$ where we can see that $a_{t}=e^{-t}$, $b_{t}=\sqrt{1-e^{-2t}}$ in Eq. (2).
@@
$\mathrm{d}y_{t}=(x_{t}+2s_{T-t}(y_{t}))\mathrm{d}t+\sqrt{2}\mathrm{d}w_{t}.$ (6)
@@ 4.2 Definitions of different solvers
where $\widehat{y}_{T-\delta}\in\mathbb{R}^{d}$ is the output, $T\in\mathbb{R}^{+}$ is the total time, and $s$ is the score estimates. The Euler-Maruyama scheme is, (Equation 5 in [CLL23]), for $t\in[t_{k}^{\prime},t_{k+1}^{\prime}]$
$\mathrm{d}\widehat{y}_{t}=(\widehat{y}_{t^{\prime}_{k}}+2s_{T-t^{\prime}_{k}}(\widehat{y}_{t^{\prime}_{k}}))\mathrm{d}t+\sqrt{2}\mathrm{d}w_{t}.$ (7)
@@
where $\widehat{y}_{T-\delta}\in\mathbb{R}^{d}$ is the output, $T\in\mathbb{R}^{+}$ is the total time, and $s$ is the score estimates. The Exponential Integrator scheme is, (Equation 6 in [CLL23]), for $t\in[t_{k}^{\prime},t_{k+1}^{\prime}]$
$\mathrm{d}\widehat{y}_{t}=(\widehat{y}_{t}+2s_{T-t^{\prime}_{k}}(\widehat{y}_{t^{\prime}_{k}}))\mathrm{d}t+\sqrt{2}\mathrm{d}w_{t}.$ (8)
@@
We define $\mathsf{DPOM}$ (Diffusion Predictor with Overdamped Modeling) as Algorithm 1 in [CCL+23] satisfying $\widehat{y}_{T-\delta}=\mathsf{DPOM}(T,h_{\mathrm{pred}},h_{\mathrm{corr}},s),$ where $\delta$ is the early stopping parameter, $\widehat{y}_{T-\delta}\in\mathbb{R}^{d}$ is the output, $T\in\mathbb{R}^{+}$ is the total steps, $h_{\mathrm{pred}}$ is the predictor step size, $h_{\mathrm{corr}}$ is the corrector step size, (see detailed definition in Algorithm 1 of [CCL+23]), and $s$ is the score estimates.
@@
We define $\mathsf{DPUM}$ (Diffusion Predictor with Underdamped Modeling) as Algorithm 2 in [CCL+23], satisfying $\widehat{y}_{T-\delta}=\mathsf{DPUM}(T,h_{\mathrm{pred}},h_{\mathrm{corr}},s),$ where $\delta$ is the early stopping parameter, $\widehat{y}_{T-\delta}\in\mathbb{R}^{d}$ is the output, $T\in\mathbb{R}^{+}$ is the total steps, $h_{\mathrm{pred}}$ is the predictor step size, $h_{\mathrm{corr}}$ is the corrector step size, (see detailed definition in Algorithm 2 of [CCL+23]), and $s$ is the score estimates.